OpenStudy (anonymous):

can you help me to solve the following inequality: x^2 >=16

OpenStudy (anonymous):

Can you show the work, please?

OpenStudy (anonymous):

graph \[y=x^2-16\] and you will see a parabola that faces up and crosses the x-axis at -4 and 4. from the picture you will see that it is positive if x<-4 or if x> 4

OpenStudy (anonymous):

analitically

OpenStudy (anonymous):

there is another more complicated way to do this, but if you know what the graph of \[y=x^2-16\] looks like it is simple.

OpenStudy (anonymous):

well step number one is going to be to write \[x^2-16\geq0\]

OpenStudy (anonymous):

so you are going to have to work with that in any case. you cannot, for example, take the square root of both sides.

OpenStudy (anonymous):

I need to do it analitically

OpenStudy (anonymous):

then you can factor \[x^2-16=(x+4)(x-4)\geq 0\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

Im such a fool

OpenStudy (anonymous):

then you can look at the sign of each factor. yes?

OpenStudy (anonymous):

you are still going to get positive, negative, positive on \[(-\infty,-4), (-4,4), (4,\infty) \] respectively

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome